# Fine-grained Attention in Hierarchical Transformers for Tabular Time-series

### v1

@@ 4.2. Models
For each dataset, we consider three hierarchical transformers: row-based, column-based and our field-based proposal $\mathsf{Fieldy}$. Additionally, we evaluate two single-stage baselines: FT-Transformer and Tabbie adapted to tabular time-series. We size all models to amount to the same total number of parameters. As $\mathsf{Fieldy}$ requires two Field transformers in its first stage, we reduce their number of layers to make the comparison fair. We keep all the other hyper-parameters related to model capacity (hidden dimensions, number of attention heads, etc.) the same across all models, similar to prior work (Padhi et al., 2021; Luetto et al., 2023; Zhang et al., 2023). Models are pre-trained for 24 epochs (Pollution) or 60 epochs (Loan default), and fine-tuned for 20 epochs. The best models are selected based on their score on a validation set and evaluated on a held-out test set.
Additionally, we include two non-deep learning baselines: XGBoost (Chen and Guestrin, 2016) and a linear model (linear or logistic regression). Note that these two baselines use exactly the same pre-processed input features and labels as the hierarchical transformers, which means that all numerical features are quantized. As for the FT-transformer variant, both of these baselines take as input a flattened version of the input tabular time-series. We select the best performing baselines after a cross-validated random search of 50 iterations to select their hyper-parameters. Also, regarding the Loan default prediction task, note that only the fine-tuning portion of the Loan dataset is considered for these non-deep learning baselines. More details on hyper-parameters can be found in Appendix B.
@@
To ensure a fair comparison across all models, we implement the same pre-training and fine-tuning strategy for all of them. While self-supervised pre-training has proved useful on tabular data (Rubachev et al., 2022), we emphasize that our objective is not to design novel pre-training techniques. Thus, we consider a simple field masking pretext task for all models, as from previous literature (Padhi et al., 2021). In detail, we guide pre-training with a BERT-like token masking pretext task (Devlin et al., 2018), randomly selecting 15% of the tokens, out of which 80% are replaced by a [MASK] token, 10% by a random token and 10% left unchanged.
Last, as our objective is not to introduce novel fine-tuning mechanisms, we resort to a standard fine-tuning technique popularized in NLP. During fine-tuning, we prepend a [CLS] token to the tabular time-series before encoding it with the model (i.e., before the Final transformer for two-stage models). Once the full time-series is encoded, this special token is extracted and passed to a final fully connected layer trained on a specific downstream task as in (Devlin et al., 2018). We implement this fine-tuning methodology for all models, with a variation for Tabbie. Indeed, in (Iida et al., 2021), the authors suggest to prepend [CLS] tokens to each row and column processed by Tabbie, based on the downstream task to learn. Thus, we consider the version of Tabbie that yields the best results in our experiments.
@@ 4.3. Results
We first remark that all transformer models outperform the non-deep learning baselines we evaluate. Although prior work demonstrated in many experiments the superiority of gradient-boosted decision trees compared to transformer-based approaches, in our setting XGBoost might suffer from using quantized numerical features. [^1] This is a consequence of our experimental protocol choices which favor comparability with existing literature (especially for transformer-related works) rather than searching for the global best independently. Note that the Pollution prediction task requires to output 20 labels for each input sample, which is implemented with a multi-output regressor wrapped on around these non-deep learning baselines, i.e., fitting one model for each target. Hence, this limits these baselines ability to capture relationships between targets. Also, for the Loan default prediction task, the performance gap is partially explained by the pre-training advantage the transformers models are given, as the non-deep learning baselines only use the smaller fine-tuning dataset.
Among transformer models, single-stage baselines underperform compared to two-stage architectures, highlighting the benefit of hierarchical representations for tabular time-series. In particular, our field-based proposal ranks first on both datasets. In contrast, TabBERT which conditions attention between fields on a unique table axis, yields worse performance indicating that capturing fields relationships across rows and columns is important. Additionally, comparing our proposal to the flattened FT-Transformer hints at the lack of table structural information for the latter. Last, while Tabbie structures the field embeddings contextualization row-wise and column-wise, its lack of a second-stage fails to relate all of these representations.
@@ 4.4. Ablation study
In Table 4, we analyze the effect of various design decisions on the performance of each transformer model on the Pollution and Loan default datasets. Note that, while the results hold qualitatively for both datasets, quantitative analysis on the Loan dataset might be affected by the smaller dataset size. In the remainder, we derive conclusions paying more attention to the Pollution dataset results.
First, we investigate the partition of model capacity between the first stage, i.e., Field transformer, and the second stage, i.e., Final transformer, for hierarchical models. Namely, we modify the number of encoder layers implemented in each stage to either favor one or the other, while ensuring the total number of model parameters remains the same (cf. Table 7). We observe that hierarchical architectures perform better when favoring the first stage. This is particularly evident for $\mathsf{Fieldy}$ (up to $+16\%$ performance improvement for the Pollution data set). This common trend indicates that the field representations learned in the first stage are particularly important for hierarchical transformers performance.
Second, we analyze the effect of the table structure encoding mechanisms that we introduced in Section 3, namely using column index embedding and/or row position embedding. Note that, even for architectures that dissect tabular time-series along table axes (i.e., Tabbie and TabBERT), the ordering of rows and columns is not preserved by design, due to the permutation-invariant nature of the basic attention mechanism, thus requiring additional positional input. Therefore, we evaluate all table structure encoding combinations for all models. We observe that explicitly indicating table structure information is beneficial to almost all transformer architectures, compared to not including any. However, the type of structural information required (i.e., row position, column index or both) is dependent on the model family.
As expected, FT-Transformer benefits from structure information on both axes. Tabbie shows a similar trend on the Loan default prediction task, although differences are less significant. Regarding two-stage models, positional encodings are particularly important for the table axis along which field representations are aggregated. Hence, column-based TabBERT benefits from row position information, conversely, row-based TabBERT exploits column index information. $\mathsf{Fieldy}$ exploits table structure information on both axes, and particularly column index embeddings.
@@ 5. Conclusion and discussion
In this paper we compared transformer architectures for tabular time-series modeling, investigating how attention can be used to simultaneously relate tabular fields across rows and columns. We evaluated our approach on tabular regression and classification tasks, showing improvements over existing baselines. However, we highlight that this work can be further expanded. In particular, we envision two possible research directions considering computational costs and more data variety and sophisticated data (pre)processing.
@@
While we ensure that in our experiments all transformer models have the same total number of parameters, their computational requirements differ. In particular, “vanilla” self-attention time complexity is $\mathcal{O}(L^{2})$ where $L$ is the input sequence length. However, these models do not consider the same input sequence, as they compute attention along various table axes, i.e., row-wise, column-wise or both. In Table 5, we compare attention complexity for each architecture. We note that the column-based and row-based hierarchical models are equivalent in terms of attention complexity. In contrast, our field-based proposal is essentially combining Tabbie-like attention in the first stage and FT-Transformer attention between all fields in the second stage. Hence, this finer-grained attention comes at a higher computational cost, which translates into longer training and inference time. Nevertheless, recent techniques for near linear-time attention (Xiong et al., 2021; Shen et al., 2021; Wang et al., 2020) can be readily used to speed up our approach.
@@
To further characterize the strengths and weaknesses of $\mathsf{Fieldy}$, we require additional evaluations on a more diverse set of tasks and datasets. In particular, the size of such datasets require careful consideration, given that over-parameterized deep learning models typically tend to overfit small datasets. In this regard, click-through rate data (Diemert Eustache, Meynet Julien et al., 2017) may be of interest for their sequential nature and large volume. [^2] We did not include it in this work because it is not considered in related literature either. Yet, given its larger size, it would be worth considering in a future work. Also, our findings have only been evaluated in the realm of tabular time-series, however, the wider domain of multi-variate time-series might also benefit from field-based hierarchical architectures, combining length and channel signals in a first stage. Finally, as the objective of this paper is limited to the comparison of attention computation axes, we do not include sophisticated tabular data embeddings techniques such as numerical features embeddings from (Gorishniy et al., 2022; Luetto et al., 2023) or target-aware pretext tasks for pre-training from (Rubachev et al., 2022). We expect their respective benefits to be portable to our novel field-based architecture, as they impact the prior embedding layer of any transformer architecture. Regarding numerical features embedding, a tangential direction may leverage methods from traditional time-series pre-processing techniques. In this regard, SAX (Lin et al., 2007) bins continuous time-series into sequences of discrete symbols to capture their trends. Such symbols could then be tokenized before being fed to a transformer-based architecture. This motivates an interesting research direction to compare numerical features embeddings schemes in the context of tabular time-series.
@@ Appendix A Field-wise attention
In this section we investigate the hypothesis that coarse-grained attention resulting from the typical hierarchical learning of tabular time-series representations might miss important cross-field relationships. More precisely, we compare TabBERT (row-based) and $\mathsf{Fieldy}$ on a simple prediction task that specifically requires the attention mechanism to focus on field dependencies across rows. For this, we utilize the Pollution dataset, sampling 100 input sequences of 10 rows each. We then [MASK] all values in the last five rows of each sample, and task the models to predict the masked tokens in the Hour column, as depicted in Figure 3. Note that values in the Hour column are always incremented by $+1$, as the Pollution measurements are hourly-based and the sequences are ordered. Hence, predicting the five missing hour values requires to specifically attend to Hour fields across separate rows.
From this simple experiment, considering the top-1 predicted token, $\mathsf{Fieldy}$ achieves an accuracy of 56%, significantly outperforming TabBERT (row-based) which only scores 9%. This indicates that the aggregation performed by hierarchical models such as TabBERT limits their ability to relate fields along their second-stage axis.
Figure 3. Illustrative input sample for the field-wise attention toy task. Predicting missing tokens in the Hour column requires field-wise attention across rows.
@@ Appendix B Hyper-parameters
Transformers architectures hyper-parameters are reported in Table 6. Note that we increase dropout and decrease hidden dimension for the Loan prediction task, as it is composed of fewer samples and models tend to overfit. We adjust the model size by selecting an appropriate number of layers as reported in Table 7. We keep all the other hyper-parameters related to model capacity the same across all models for a fair comparison.

### v2

@@ 4.2. Models
For each dataset, we consider three hierarchical transformers: row-based, column-based and our field-based proposal $\mathsf{Fieldy}$. We use the official code from TabBERT (Padhi et al., 2021) to implement row-based and column-based baselines. Additionally, we evaluate two single-stage baselines: FT-Transformer and Tabbie, using our own implementation to adapt them to tabular time-series. We size all models to amount to the same total number of parameters. As $\mathsf{Fieldy}$ requires two Field transformers in its first stage, we reduce their number of layers to make the comparison fair. We keep all the other hyper-parameters related to model capacity (hidden dimensions, number of attention heads, etc.) the same across all models, similar to prior work (Padhi et al., 2021; Luetto et al., 2023; Zhang et al., 2023). Models are pre-trained for 24 epochs (Pollution) or 60 epochs (Loan default), and fine-tuned for 20 epochs. The best models are selected based on their score on a validation set and evaluated on a held-out test set. Additionally, we include two non-deep learning baselines: XGBoost (Chen and Guestrin, 2016) and a linear model (linear or logistic regression). Note that these two baselines use exactly the same pre-processed input features and labels as the hierarchical transformers, which means that all numerical features are quantized. Both of these baselines take as input a flattened version of the tabular time-series. We select the best linear models after a cross-validated random search of 50 iterations to select their hyper-parameters. Also, regarding the Loan default prediction task, note that only the fine-tuning portion of the Loan dataset is considered for these non-deep learning baselines. More details on hyper-parameters can be found in Appendix B.
@@
To ensure a fair comparison across all models, we implement the same pre-training and fine-tuning strategy for all of them. While self-supervised pre-training has proved useful on tabular data (Rubachev et al., 2022), we emphasize that our objective is not to design novel pre-training techniques. Thus, we consider a simple field masking pretext task for all models, as from previous literature (Padhi et al., 2021). In detail, we guide pre-training with a BERT-like token masking pretext task (Devlin et al., 2018), randomly selecting 15% of the tokens, out of which 80% are replaced by a [MASK] token, 10% by a random token and 10% left unchanged. Last, as our objective is not to introduce novel fine-tuning mechanisms, we resort to a standard fine-tuning technique popularized in NLP. During fine-tuning, we prepend a [CLS] token to the tabular time-series before encoding it with the model (i.e., before the Final transformer for two-stage models). Once the full time-series is encoded, this special token is extracted and passed to a final fully connected layer trained on a specific downstream task as in (Devlin et al., 2018). We implement this fine-tuning methodology for all models, with a variation for Tabbie. Indeed, in (Iida et al., 2021), the authors suggest to prepend [CLS] tokens to each row and column processed by Tabbie, based on the downstream task to learn. Thus, we consider the version of Tabbie that yields the best results in our experiments.
@@ 4.3. Results
We first remark that all transformer models outperform the non-deep learning baselines we evaluate. [^1] Although prior work demonstrated in many experiments the superiority of gradient-boosted decision trees compared to transformer-based approaches, in our setting, XGBoost might suffer from using quantized numerical features. This is a consequence of our experimental protocol choices which favor comparability with existing literature (especially for transformer-related works) rather than searching for the global best independently. Note that the Pollution prediction task requires to output 20 labels for each input sample, which is implemented with a multi-output regressor wrapped on around these non-deep learning baselines, i.e., fitting one model for each target. Hence, this limits these baselines’ ability to capture relationships between targets. Also, for the Loan default prediction task, the performance gap is partially explained by the pre-training advantage the transformers models are given, as the non-deep learning baselines only use the smaller fine-tuning dataset.
Among transformer models, single-stage baselines underperform compared to two-stage architectures, highlighting the benefit of hierarchical representations for tabular time-series. In particular, our field-based proposal ranks first on both datasets. In contrast, TabBERT which conditions attention between fields on a unique table axis, yields worse performance indicating that capturing fields relationships across rows and columns is important. Additionally, comparing $\mathsf{Fieldy}$ to the flattened FT-Transformer hints at the lack of table structural information for the latter. Last, while Tabbie structures the field embeddings contextualization row-wise and column-wise, its lack of a second-stage fails to relate all of these representations.
@@ 4.4. Ablation study
In Table 4, we analyze the effect of various design decisions on the performance of each transformer model on the Pollution and Loan default datasets. Note that, while the results hold qualitatively for both datasets, quantitative analysis on the Loan task might be affected by the smaller dataset size. In the remainder, we derive conclusions paying more attention to the Pollution dataset results.
First, we investigate the partition of model capacity between the first stage, i.e., Field transformer, and the second stage, i.e., Final transformer, for hierarchical models. Namely, we modify the number of encoder layers implemented in each stage to either favor one or the other, while ensuring the total number of model parameters remains the same (cf. Table 7). We observe that hierarchical architectures perform better when favoring the first stage. This is particularly evident for $\mathsf{Fieldy}$ with up to $+16\%$ performance improvement for the Pollution data set. This common trend indicates that the field representations learned in the first stage are particularly important for hierarchical transformers performance.
Second, we analyze the effect of the table structure encoding mechanisms that we introduced in Section 3, namely using column index embedding and/or row position embedding. Note that, even for architectures that dissect tabular time-series along table axes (i.e., Tabbie and TabBERT), the ordering of rows and columns is not preserved by design, due to the permutation-invariant nature of the basic attention mechanism, thus requiring additional positional input. Therefore, we evaluate all table structure encoding combinations for all models. We observe that explicitly indicating table structure information is beneficial to almost all transformer architectures, compared to not including any. However, the type of structural information required (i.e., column index, row position, or both) is dependent on the model family. As expected, FT-Transformer benefits from structure information on both axes. Tabbie shows a similar trend on the Loan default prediction task, although differences are less significant. Regarding two-stage models, positional encodings are particularly important for the table axis along which field representations are aggregated. Hence, column-based TabBERT benefits from row position information, conversely, row-based TabBERT exploits column index information. $\mathsf{Fieldy}$ exploits table structure information on both axes, and particularly column index embeddings.
@@ 5. Conclusion and discussion
In this paper we compared transformer architectures for tabular time-series modeling, investigating how attention can be used to simultaneously relate tabular fields across rows and columns. We evaluated our approach on tabular regression and classification tasks, showing improvements over existing baselines. However, we highlight that this work can be further expanded. In particular, we envision two possible research directions considering computational costs and more data variety.
@@
While we ensure that in our experiments all transformer models have the same total number of parameters, their computational requirements differ. In particular, “vanilla” self-attention time complexity is $\mathcal{O}(L^{2})$ where $L$ is the input sequence length. However, the transformer models we compare do not consider the same input sequence, as they compute attention along various table axes, i.e., row-wise, column-wise or both. In Table 5, we compare attention complexity for each architecture, based on the number of rows and columns composing the tabular time-series. We note that the column-based and row-based hierarchical models are equivalent in terms of attention complexity. In contrast, our field-based proposal is essentially combining Tabbie-like attention in the first stage and FT-Transformer attention between all fields in the second stage. Hence, this finer-grained attention comes at a higher computational cost, which translates into longer training and inference time. Nevertheless, recent techniques for near linear-time attention (Xiong et al., 2021; Shen et al., 2021; Wang et al., 2020) can be readily used to speed up our approach.
@@
To further characterize the strengths and weaknesses of $\mathsf{Fieldy}$, we require additional evaluations on a more diverse set of tasks and datasets. In particular, the size of such datasets require careful consideration, given that over-parameterized deep learning models typically tend to overfit small datasets. In this regard, click-through rate data (Diemert Eustache, Meynet Julien et al., 2017) may be of interest for their sequential nature and large volume. We did not include such dataset in this work because it is not considered in related literature either. Yet, given its larger size, it would be worth considering in a future work. Also, our findings have only been evaluated in the realm of tabular time-series, however, the wider domain of multi-variate time-series might also benefit from field-based hierarchical architectures, combining length and channel signals in a first stage. Finally, as the objective of this paper is limited to the comparison of attention computation axes, we do not include sophisticated tabular data embeddings techniques such as numerical features embeddings from (Gorishniy et al., 2022; Luetto et al., 2023) or target-aware pretext tasks for pre-training from (Rubachev et al., 2022). We expect their respective benefits to be portable to our novel field-based architecture, as they impact the prior embedding layer of any transformer architecture. Regarding numerical features embedding, a tangential direction may leverage methods from traditional time-series pre-processing techniques. In this regard, SAX (Lin et al., 2007) bins continuous time-series into sequences of discrete symbols to capture their trends. Such symbols could then be tokenized before being fed to a transformer-based architecture. This motivates an interesting research direction to compare numerical features embeddings schemes in the context of tabular time-series.
@@ Appendix A Field-wise attention
In this section, we investigate the hypothesis that coarse-grained attention resulting from the typical hierarchical learning of tabular time-series representations might miss important cross-field relationships. More precisely, we compare a pre-trained TabBERT (row-based architecture) and a pre-trained $\mathsf{Fieldy}$ (field-based architecture) on a simple prediction task that specifically requires the attention mechanism to focus on field dependencies across rows. For this, we utilize the Pollution dataset introduced in Section 4, sampling 100 input sequences of 10 rows each. We then [MASK] all values in the last five rows of each sample, and task the models to predict the masked tokens in the Hour column, as depicted in Figure 3. Note that values in the Hour column are always incremented by $+1$, as the Pollution measurements are hourly-based and the sequences are ordered. Hence, predicting the five missing hour values requires to specifically attend to Hour fields across separate rows. From this simple experiment, considering the top-1 predicted token, $\mathsf{Fieldy}$ (field-based) achieves an accuracy of 56%, significantly outperforming TabBERT (row-based) which only scores 9%. This indicates that the aggregation performed by hierarchical models such as TabBERT limits their ability to relate fields along their second-stage axis.
@@ Appendix B Hyper-parameters
Transformers architectures hyper-parameters are reported in Table 6. Note that we increase dropout and decrease the hidden dimension for the Loan prediction task, as it is composed of fewer samples and models tend to overfit. We adjust the model size by selecting an appropriate number of layers as reported in Table 7. We keep all the other hyper-parameters related to model capacity the same across all models for a fair comparison.